# Conditional Density Estimation with Neural Networks for Bias Correction of Multivariate Climate Model Data

### v1

@@ 2.7 Model evaluation metrics for SPCDE
In the study presented in Section 3, we assess the efficacy of SPCDE based on five functionals of the joint distribution of TMAX and PRCP: the Wasserstein distance between the marginal densities of calibrated and observational data, lag-1 autocorrelations, pairwise spatial correlations, and their 0.95 quantiles. These are evaluated for both the observed (nClimGrid) and bias-corrected predictions. We also evaluated the proportion of zeros for PRCP, and the cross correlation between TMAX and PRCP. The Wasserstein distance (Kantorovich, 1960; Vasershtein, 1969) quantifies the overall fit of the marginal distributions of each variable in terms of the dissimilarity between the densities of the observed and bias-corrected data, whereas the proportion of zeros and the 0.95 quantiles focus specifically on the tail behavior of the distributions. The three correlation metrics ensure that the bias-corrected data maintains the dependence structure of the nClimGrid data.
In addition to assessing goodness of fit of our bias corrected data, we also compare it with QM and CCA. As described in Section 1, QM is a bias correction method which matches the quantiles of the observational data time series and the GCM time series. The asynchronous QM bias correction method sorts GCM and nClimGrid data vectors in ascending order (independently), and then fits linear regression to the sorted data. This can take the form of simple linear regression (e.g., Dettinger et al., 2004) or piecewise linear regression (Stoner et al., 2013). The QM methodology is applied separately for each month, grid cell, and variable (TMAX and PRCP). The goal is to obtain a linear map between the quantiles of the GCM and nClimGrid datasets. We can then provide GCM data for unobserved grid points or time periods as input to the fitted model and obtain bias-corrected TMAX and PRCP as output. The resulting values must subsequently be rearranged back into the original order to retrieve the final product, a continuous chronological time series of the bias-corrected values. Since QM in our example is applied to each variable separately, it fails to account for inter-variable dependencies (Bhowmik and Sankarasubramanian, 2019). This can distort relationships critical for understanding processes such as hydrological cycles (Seo et al., 2019). We use CCA (Bhowmik et al., 2017) as our other point of comparison. CCA employs multivariate sorting of variables based on their joint probability of occurrence such that their cross correlation is maintained. Afterwards, canonical correlation is used to bias correct variables simultaneously. Neither of these two approaches, however, explicitly models spatial or temporal dependence present in the data.
3 Bias correction of historical TMAX and PRCP
@@ 3.1 Overview
We implemented our approach for daily observational (nClimGrid) and model (CMIP6) data from 1951–2014 on two $5\times 5$ grids as shown in Figure 1. We separated the data into a training period (1951–2000) and a testing period (2001–2014), and fitted the spatiotemporal SPQR bias correction model to the training period data. The Vecchia neighbor sets are selected using max-min ordering, which can result in significant improvements over other coordinate based orderings (Katzfuss and Guinness, 2021). We model each month’s data separately to account for seasonality. Additionally, since daily PRCP contains a large proportion of zeroes and SPQR does not work with semi-continuous data, we do not model PRCP directly. Instead, we model $\log(0.0001+\mbox{PRCP})$, a common approach when modeling precipitation data using continuous distributions (Rajagopalan and Lall, 1998). We found this transformation sufficient for the simultaneous estimation of the proportion of zeros in the data, combined with a post-processing step described below, as well as high quantiles.
@@
The SPQR models (for each month, location, and variable) can be fit in parallel, since the Vecchia approximation decomposes the joint density of Y into independent univariate conditional densities. Predictions for every month can also be carried out independently, and therefore the calibration procedure occurs in parallel. Within each month, however, predictions need to be made in the same sequence as the ordering of the Vecchia neighbor sets. Predictions for TMAX and PRCP were obtained for the testing period of 2001–2014 at each location using Algorithm 2.
@@ 3.2 Evaluating bias-corrected output based on SPCDE
We first present results from bias correction at both sets of locations for the validation period of 2001–2014. Figure 2 plots density curves of bias corrected TMAX and log-scale PRCP for the SE and SW regions. Each curve pools the entire validation period data across all 25 locations, and we see that the calibrated TMAX and PRCP closely match the observed data. Note that the density plots for PRCP are displayed in the transformed log scale since visual inspection of the density plots are most informative in the scale that the data was analyzed. However, all remaining plots and metrics for PRCP are presented in the raw data scale. Fitting separate monthly models for TMAX captures the pronounced bimodal distribution in the SW region with different modes for observed and model data. On the other hand, there are substantial differences between the observed and model densities of PRCP at both tails of their respective distributions. For both variables, SPQR is able to successfully calibrate the model data to match the density of the observational data.
@@
Figures 34 plot the 0.95 quantiles of TMAX and PRCP, and the proportion of zeros for PRCP, respectively. Each panel displays the effect of calibrating the model data by comparing it against the observational data for every location and month. For the upper 0.95 quantiles, the uncalibrated model values for TMAX have noticeable bias, something that can be visually verified in Figure 2. The calibrated quantiles virtually eliminates this bias while also having lower uncertainty than the uncalibrated data. The calibration also improves upper quantile estimation for PRCP. However, it is unable to fully adjust for the bias, and there is noticeably more uncertainty in the PRCP estimates than TMAX. Part of this bias is attributable to the proportion of zeros in PRCP data, which varies significantly by season. We therefore treat model precipitation below $0.001$ mm as zero precipitation, since it is not measurable by physical instruments like rain gauges. This was also done for the SPQR estimates since the continuous fitted distribution for PRCP will never produce exact zeros. The uncalibrated model data (red) in Figure 4 severely underestimates the proportion of zeros, due to a well known phenomenon called the drizzle effect (see, e.g., Levey and Sankarasubramanian, 2024), where numerical climate models for precipitation very rarely generate zero precipitation values, instead usually having very small but positive values. The calibrated density of PRCP is able to accurately represent the proportion of zeros that are observed in the real data, indicating that our log-transform for modeling PRCP is adequate for this application.
@@
Figures 56 plot the effect of calibration on the autocorrelations for each variable, and the cross correlation between them, respectively. As in the previous two plots, points correspond to month and location specific estimates. In Figure 5, the calibration is unable to improve the autocorrelation for SW TMAX from its model values, but for the remaining three variables, the calibrated data has lower bias and variability in its estimates of autocorrelation compared to the uncalibrated model data. For the cross correlations in Figure 6, the uncalibrated data has positive bias in the cross correlations for both regions. The calibration was able to virtually eliminate the bias and provided highly accurate and precise estimates of the cross correlation. This is in part because the conditional density of PRCP depends on both current and past TMAX values, and therefore SPQR implicitly models the dependence between the two variables.
@@
Finally, Figure 7 shows pairwise spatial correlations for the calibrated and uncalibrated data. Each panel contains values for 300 pairs of locations. Values are concentrated around the diagonal, suggesting the the calibrated data has similar spatial correlation compared to the observed data. We do notice more uncertainty in the lower end of TMAX correlation values in the uncalibrated data, that the calibration is not able to entirely address, but overall the calibrated data has lower bias and uncertainty compared to the uncalibrated model data.
3.3 Comparison of SPCDE with competing methods
To evaluate the performance of the SPCDE bias correction approach compared to CCA and QM, we provide metrics for predictive accuracy based on the three methods in Table 1. Marginal fits for TMAX and the transformed PRCP are summarized using the Wasserstein distance, while for the remainder of the metrics provided in this section, we evaluate the RMSE. In each case, smaller values indicate more accurate predictions. We note that the SPCDE predictions are better than CCA and QM predictions across the board. The only two instances where SPCDE is not outright better are the 0.95 quantiles for SW TMAX and PRCP. Neither CCA nor QM is clearly better even in this case, with the former having lower RMSE for TMAX, and the latter having a lower RMSE for PRCP. On the other hand, SPCDE is still the best among the three methods at estimating the proportion of zeros in PRCP for both regions. The autocorrelations are also significantly better for SPCDE, since neither of the competing method explicitly accounts for autocorrelation. While SPCDE has smaller RMSE for spatial correlations in every case, the differences are more noticeable for the SE, which has both lower bias and variability in its spatial correlation estimates (see Figure 7). Finally, RMSE in the cross correlations are around 3 three orders of magnitude smaller for SPCDE than competing methods.
Table 1: Out-of-sample prediction accuracy in the bias correction of TMAX and PRCP based on a spatiotemporal SPCDE model, compared with canonical correlation analysis (CCA) and quantile mapping (QM). Other than the Wasserstein distance metric, the remaining values correspond to the RMSE between observed and calibrated data for 2001–2014. Metrics are averaged across spatial locations (pairs of locations for spatial correlation). Best (smallest) values for each combination of region, variable, and metric are highlighted.
@@ 4 Discussion
In this paper, we develop a new SPCDE approach for the simultaneous bias correction of spatiotemporal PRCP and TMAX. A key goal of our study is to maintain the cross correlation between the two variables during bias correction. SPCDE leverages the Vecchia approximation to maintain spatiotemporal and inter-variable dependencies, and density estimation is carried out using the neural network based SPQR method. Our approach is trained across two spatial fields in the Southeast and Southwest US using historical data from 1951–2000, and tested using data from 2001–2014 for the same region. SPCDE is able to outperform QM and CCA, two competing methods, and the joint distribution characteristics of the observational data is preserved in the bias-corrected data.
@@ Appendix A Additional results
Table 2: Out-of-sample prediction accuracy in the bias correction of TMAX and PRCP based on a spatiotemporal SPQR model, compared with canonical correlation analysis (CCA) and quantile mapping (QM). Other than the Wasserstein distance metric, the remaining values correspond to the RMSE between observed and calibrated data for 1951–2014. Metrics are averaged across spatial locations (pairs of locations for spatial correlation). Best (smallest) values for each combination of region, variable, and metric are highlighted.

### v2

@@ 2.7 Model evaluation metrics for SPCDE
In the study presented in Section 3, we assess the efficacy of SPCDE based on five functionals of the joint distribution of TMAX and PRCP: the Wasserstein distance between the marginal densities of calibrated and observational data, lag-1 autocorrelations, pairwise spatial correlations, and their 0.95 quantiles. These are evaluated for both the observed (nClimGrid) and calibrated predictions. We also evaluated the proportion of zeros for PRCP, and the cross correlation between TMAX and PRCP. The Wasserstein distance (Kantorovich, 1960; Vasershtein, 1969) quantifies the overall fit of the marginal distributions of each variable in terms of the dissimilarity between the densities of the observed and calibrated data, whereas the proportion of zeros and the 0.95 quantiles focus specifically on the tail behavior of the distributions. The three correlation metrics ensure that the calibrated data maintains the dependence structure of the nClimGrid data.
In addition to assessing goodness of fit of our density corrected data, we also compare it with QM and CCA. As described in Section 1, QM is a bias correction method which matches the quantiles of the observational data time series and the GCM time series. The asynchronous QM bias correction method sorts GCM and nClimGrid data vectors in ascending order (independently), and then fits linear regression to the sorted data. This can take the form of simple linear regression (e.g., Dettinger et al., 2004) or piecewise linear regression (Stoner et al., 2013). The QM methodology is applied separately for each month, grid cell, and variable (TMAX and PRCP). The goal is to obtain a linear map between the quantiles of the GCM and nClimGrid datasets. We can then provide GCM data for unobserved grid points or time periods as input to the fitted model and obtain bias-corrected TMAX and PRCP as output. The resulting values must subsequently be rearranged back into the original order to retrieve the final product, a continuous chronological time series of the bias-corrected values. Since QM in our example is applied to each variable separately, it fails to account for inter-variable dependencies (Bhowmik and Sankarasubramanian, 2019). This can distort relationships critical for understanding processes such as hydrological cycles (Seo et al., 2019). We use CCA (Bhowmik et al., 2017) as our other point of comparison. CCA employs multivariate sorting of variables based on their joint probability of occurrence such that their cross correlation is maintained. Afterwards, canonical correlation is used to bias correct variables simultaneously. Neither of these two approaches, however, explicitly models spatial or temporal dependence present in the data.
3 Density correction of historical TMAX and PRCP
@@ 3.1 Overview
We implemented SPCDE for daily observational and GCM data from 1951–2014 on two separate $5\times 5$ grids as shown in Figure 1. We separated the data into a training period (1951–2000) and a testing period (2001–2014), and fitted the spatiotemporal SPCDE model to the training period data. The Vecchia neighbor sets are selected using max-min ordering, which can result in significant improvements over other coordinate based orderings (Guinness, 2018). We model each month’s data separately to account for seasonality. Additionally, since daily PRCP contains a large proportion of zeroes and SPQR does not work with semi-continuous data, we do not model PRCP directly. Instead, we model $\log(0.0001+\mbox{PRCP})$, a common approach when modeling precipitation data using continuous distributions (Rajagopalan and Lall, 1998). We found this transformation sufficient for the simultaneous estimation of the proportion of zeros in the data, combined with a post-processing step (described in Section 3.2), as well as high quantiles.
@@
The SPQR models (for each month, location, and variable) can be fit in parallel, since the Vecchia approximation decomposes the joint density of Y into independent univariate conditional densities. Predictions for every month can also be carried out independently, and therefore the projection/calibration step occurs in parallel. Within each month, however, predictions need to be made in the same sequence as the ordering of the Vecchia neighbor sets. Predictions for TMAX and PRCP were obtained for the testing period of 2001–2014 at each location using Algorithm 2.
@@ 3.2 Evaluating bias-corrected output based on SPCDE
We first present results from density correction at both sets of locations for the validation period of 2001–2014. Figure 2 plots density curves of density corrected TMAX and log-scale PRCP for the SE and SW regions. Each curve pools the entire validation period data across all 25 locations, and we see that the calibrated TMAX and PRCP closely match the observed data. Note that the density plots for PRCP are displayed in the transformed log scale since visual inspection of the density plots are most informative in the scale that the data was analyzed. However, all remaining plots and metrics for PRCP are presented in the raw data scale. Fitting separate monthly models for TMAX captures the pronounced bimodal distribution in the SW region with different modes for observed and model data. On the other hand, there are substantial differences between the observed and model densities of PRCP at both tails of their respective distributions. For both variables, SPQR is able to successfully calibrate the model data to match the density of the observational data.
@@
Figures 34 plot the 0.95 quantiles of TMAX and PRCP, and the proportion of zeros for PRCP, respectively. Each panel displays the effect of density correction by comparing it against the observational data for every location and month. For the upper 0.95 quantiles, the uncalibrated model values for TMAX have noticeable bias, something that can be visually verified in Figure 2. The calibrated quantiles virtually eliminates this bias while also having lower uncertainty than the uncalibrated data. The density correction also improves upper quantile estimation for PRCP. However, it is unable to fully adjust for the bias, and there is noticeably more uncertainty in the PRCP estimates than TMAX. Part of this bias is attributable to the proportion of zeros in PRCP data, which varies significantly by season. We therefore treat model precipitation below $0.001$ mm as zero precipitation, since it is not measurable by physical instruments like rain gauges. This was also done for the SPQR estimates since the continuous fitted distribution for PRCP will never produce exact zeros. The uncalibrated model data (red) in Figure 4 severely underestimates the proportion of zeros, due to a well known phenomenon called the drizzle effect (see, e.g., Levey and Sankarasubramanian, 2024), where numerical climate models for precipitation very rarely generate zero precipitation values, instead usually having very small but positive values. The calibrated density of PRCP is able to accurately represent the proportion of zeros that are observed in the real data, indicating that our log-transform for modeling PRCP is adequate for this application.
@@
Figures 56 plot the effect of density correction on the autocorrelations for each variable, and the cross correlation between them, respectively. As in the previous two plots, points correspond to month and location specific estimates. In Figure 5, SPCDE is unable to improve the autocorrelation for SW TMAX from its model values, but for the remaining three variables, the calibrated data has lower bias and variability in its estimates of autocorrelation compared to the uncalibrated model data. For the cross correlations in Figure 6, the uncalibrated data has positive bias in the cross correlations for both regions. Density correction was able to virtually eliminate the bias and provided highly accurate and precise estimates of the cross correlation. This is in part because the conditional density of PRCP depends on both current and past TMAX values, and therefore SPQR models the dependence between the two variables.
Finally, Figure 7 shows pairwise spatial correlations for the calibrated and uncalibrated data. Each panel contains values for 300 pairs of locations. Values are concentrated around the diagonal, suggesting the the calibrated data has similar spatial correlation compared to the observed data. We do notice more uncertainty in the lower end of TMAX correlation values in the uncalibrated data, that the density correction is not able to entirely address, but overall the calibrated data has lower bias and uncertainty compared to the uncalibrated model data.
3.3 Comparison of SPCDE with CCA and QM
Table 1: Out-of-sample prediction accuracy in the density correction of TMAX and PRCP based on a spatiotemporal SPCDE model, compared with canonical correlation analysis (CCA) and quantile mapping (QM). Other than the Wasserstein distance metric, the remaining values correspond to the RMSE between observed and calibrated data for 2001–2014. Metrics are averaged across spatial locations (pairs of locations for spatial correlation). Best (smallest) values for each combination of region, variable, and metric are highlighted.
@@ 3.3 Comparison of SPCDE with CCA and QM
To evaluate the performance of the SPCDE density correction approach compared to CCA and QM, we provide metrics for predictive accuracy based on the three methods in Table 1. Marginal fits for TMAX and the transformed PRCP are summarized using the Wasserstein distance, while for the remainder of the metrics provided in this section, we evaluate the RMSE. In each case, smaller values indicate more accurate predictions. We note that the SPCDE predictions are better than CCA and QM predictions across the board. The only two instances where SPCDE is not outright better are the 0.95 quantiles for SW TMAX and PRCP. Neither CCA nor QM is clearly better even in this case, with the former having lower RMSE for TMAX, and the latter having a lower RMSE for PRCP. On the other hand, SPCDE is still the best among the three methods at estimating the proportion of zeros in PRCP for both regions. The autocorrelations are also significantly better for SPCDE, since neither of the competing method explicitly accounts for autocorrelation. While SPCDE has smaller RMSE for spatial correlations in every case, the differences are more noticeable for the SE, which has both lower bias and variability in its spatial correlation estimates (see Figure 7). Finally, RMSE in the cross correlations are around 3 three orders of magnitude smaller for SPCDE than competing methods.
@@ 4 Discussion
In this paper, we develop a new SPCDE approach for the simultaneous density correction of spatiotemporal PRCP and TMAX. A key goal of our study is to maintain the cross correlation between the two variables during density correction. SPCDE leverages the Vecchia approximation to maintain spatiotemporal and inter-variable dependencies, and density estimation is carried out using the deep learning based SPQR method. Our approach is trained across two spatial fields in the Southeast and Southwest US using historical data from 1951–2000, and tested using data from 2001–2014 for the same region. SPCDE is able to outperform QM and CCA, two competing methods, and the joint distribution characteristics of the observational data is preserved in the bias-corrected data.
@@ Appendix A Additional results
Table 2: Out-of-sample prediction accuracy in the density correction of TMAX and PRCP based on a spatiotemporal SPQR model, compared with canonical correlation analysis (CCA) and quantile mapping (QM). Other than the Wasserstein distance metric, the remaining values correspond to the RMSE between observed and calibrated data for 1951–2014. Metrics are averaged across spatial locations (pairs of locations for spatial correlation). Best (smallest) values for each combination of region, variable, and metric are highlighted.